Suppose you buy an appliance costing ​$1650 at a store charging 14​% ​add-on interest, you make a down payment of $ 800​, and you agree to monthly payments over 4 years. Find the total cost for the appliance plus interest.
Answer by Boreal(15235)

payoff $850
m=r(PV)/1-(1+r)^(-n); rPV=(.14/12)(850)=9.9166667
1-(1+r)^(-n)=0.4269
divide unrounded 9.92/0.4269 to get $23.23